OpenStudy (anonymous):

Determine the first five iterates of f(x) = 1.5x(1 – x) for x0 = 0.8. (1 point) a. 0.245, 0.25, 0.255, 0.26, 0.265 b. 0.243, 0.246, 0.249, 0.252 c. 1.24, 2.24, 3.24, 4.24, 5.24 d. 0.24, 0.274, 0.298, 0.314, 0.323

OpenStudy (anonymous):

Firstly, do you understand the question?

OpenStudy (anonymous):

It means f^5(0.8) = f(f(f(f(f(0.8))))

OpenStudy (anonymous):

Not to much its my 2nd day into this subject

OpenStudy (anonymous):

f(x) = 1.5x(1 - x) f(0.8) = 1.5(0.8)(1 - 0.8) = 0.24 f(0.24) = 1.5(0.24)(1 - 0.24) ~= 0.274 f(0.274) = 1.5(0.274)(1 - 0.274) ~= 0.298 f(0.298) = 1.5(0.298)(1 - 0.298) ~= 0.314 f(0.314) = 1.5(0.314)(1 - 0.314) ~= 0.323

OpenStudy (anonymous):

So the answer is D

OpenStudy (anonymous):

thanks im trying to get a hang on this on just started